OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (bahrom7893):

|2x+6| -4 + 4 = 20 + 4 |2x+6| = 24

OpenStudy (bahrom7893):

Which means that 2x+6 could be either positive or negative 24, because: |24| = 24 and |-24|=24, so: (2x+6) = 24 or -(2x+6) = 24

OpenStudy (bahrom7893):

2x + 6 = 24 2x = 24-6 = 18 x=9

OpenStudy (excalibur0126):

It's C but bahrom7893 can explain it

OpenStudy (bahrom7893):

-(2x+6) = 24 2x+6 = -24 2x = -24-6 = -30 x = -30/2 = -15

OpenStudy (anonymous):

oooh ok i see thank you :) can u help me with one more?

OpenStudy (anonymous):

@bahrom7893

hero (hero):

\(|2x + 6| = 24\) \(\sqrt{(2x + 6)^2}= 24\) \((2x + 6)^2 = 24^2\) \(4x^2 + 24x + 36 = 576\) \(4x^2 + 24x - 540 = 0\) \(4(x^2 + 6x - 135 = 0\) \(x^2 + 6x - 135 = 0\) \(x^2 + (15 - 9)x - 135 = 0\) \(x^2 + 15x - 9x - 135 = 0\) \(x(x + 15) - 9(x + 15)\) \((x + 15)(x - 9) = 0\) \(x = \{-15, 9\}\)